# Functional Stochastic Gradient MCMC for Bayesian Neural Networks

### v1

@@ Abstract
Classical variational inference for Bayesian neural networks (BNNs) in parameter space usually suffers from unresolved prior issues such as knowledge encoding intractability and pathological behaviors in deep networks, which could lead to an improper posterior inference. Hence, functional variational inference has been proposed recently to resolve these issues via stochastic process priors. Beyond variational inference, stochastic gradient Markov Chain Monte Carlo (SGMCMC) is another scalable and effective inference method for BNNs to asymptotically generate samples from true posterior by simulating a continuous dynamic. However, the existing SGMCMC methods only work in parametric space, which has the same issues of parameter-space variational inference, and extending the parameter-space dynamics to function-space dynamics is not a trivial undertaking. In this paper, we introduce a new functional SGMCMC scheme via newly designed diffusion dynamics, which can incorporate more informative functional priors. Moreover, we prove that the stationary distribution of these functional dynamics is the target posterior distribution over functions. We demonstrate better performance in both accuracy and uncertainty quantification of our functional SGMCMC on several tasks compared with naive SGMCMC and functional variational inference methods.
Machine Learning, ICML
@@ 1 Introduction
Compared to regular deep neural networks that rely on a single pattern of model parameters only, Bayesian neural networks (BNNs) have demonstrated the ability to provide a principled predictive uncertainty estimation and enhanced generalization performance through a model integration under posterior distribution over network weights (Neal, 1995; Gal et al., 2016; Wilson & Izmailov, 2020). With the continued development and success of Bayesian inference methods in the deep learning community, BNNs have been applied in a wide variety of domains and tasks, including prediction in astrophysics (Cranmer et al., 2021), medical diagnosis and diabetic detection (Filos et al., 2019; Band et al., 2022), and other safety-critical environments (Rudner et al., 2022).
@@
Prior choosing is apparently crucial for BNNs. The existing common practice is i.i.d. Gaussian priors for all network weights, which is also named parameter-space prior. However, this parameter-space prior has been shown to have some problematic issues. For example, as network depth increases, the function samples of Gaussian prior over weights tend to be horizontal (Duvenaud et al., 2014; Matthews et al., 2018), and the prior distribution of unit outputs becomes more heavy-tailed with ReLU activation (Vladimirova et al., 2019; Tran et al., 2022), which will lead to improper posterior inference (Papamarkou et al., 2022). Moreover, the effects of this parameter space prior to the posterior inference remain an open question due to the complex network architecture (Fortuin et al., 2021; Wild et al., 2022), so it is difficult to transform the valid prior knowledge into corresponding network weights’ prior. Given these unresolved prior problems for parameters, Bayesian inference directly in function space has gained increasing interest in recent years since one can incorporate more informative stochastic processes as the prior. For example, Gaussian processes (GPs) which can easily encode the prior knowledge about the potential functions via kernel functions choosing have been widely used as the prior in function-space variational inference for BNNs (Sun et al., 2019; Ma & Hernández-Lobato, 2021; Rudner et al., 2022; Pielok et al., 2022) which has been shown to yield better prediction and improved uncertainty estimation performance than variational inference in parameter space.
@@
For the posterior inference given prior, Markov Chain Monte Carlo (MCMC) is another valid method for BNNs apart from the commonly used variational inference (Blundell et al., 2015; Sun et al., 2019). Instead of optimizing a variational posterior by minimizing the distance (divergence) with the true posterior (Graves, 2011; Blundell et al., 2015; Gal & Ghahramani, 2016), MCMC approaches are regarded as the gold standard for Bayesian inference since they generate discrete samples from a continuous-time Markov chain whose stationary distribution is exactly the true posterior (Neal et al., 2011; Welling & Teh, 2011; Li et al., 2016). So, they do not need to make strong assumptions about the distributional form of the posterior (Alexos et al., 2022; Cobb & Jalaian, 2021). Even better, the computationally expensive MCMC methods became feasible for modern BNNs with large data following the introduction of their scalable stochastic gradient variants (SGMCMC) (Welling & Teh, 2011; Patterson & Teh, 2013; Chen et al., 2014; Zhang et al., 2020; Nemeth & Fearnhead, 2021). However, all existing MCMC methods for BNNs are still in parameter space, and extending them to function space is not a minor process because we need to design novel dynamics in function space and ensure the new dynamics would have the target posterior as the stationary distribution.
@@
In this work, we propose a functional SGMCMC scheme based on the fundamental Itô lemma for stochastic calculus (Itô, 1951; Brzeźniak et al., 2008; Øksendal, 2013). The dynamics of BNNs in function space are designed by proposing a new potential energy function where we integrate the more informative functional prior, such as GPs, into the diffusion process. Furthermore, we prove that the stationary distribution of the new functional dynamics is the desired posterior distribution over functions. Our main contributions are as follows:
  • We propose a new functional SGMCMC approach for BNNs that can utilize the more informative functional prior to the posterior inference.
@@
  • We prove that under our designed functional SGMCMC dynamics, the stationary distribution would be the target posterior distribution over functions of BNNs.
@@
  • Our method shows improved predictive performance and uncertainty estimation in several benchmark tasks compared to parameter-space SGMCMC and parameter- and function-space variational inference methods.
@@ 2.1 Bayesian Neural Networks
[…]
@@ 2.2 Stochastic Gradient MCMC for BNNs
Traditional MCMC methods, such as the random-walk Metropolis-Hastings (MH) algorithm (Metropolis et al., 1953), are easy to apply but suffer from a slow mixing rate and high correlation between samples. In contrast, the dynamics-based sampling methods utilizing gradients of log posterior density can provide a more efficient exploration of the posterior (Girolami & Calderhead, 2011). Also, their stochastic gradient variants (Welling & Teh, 2011; Chen et al., 2014; Baker et al., 2019), such as the stochastic gradient Langevin dynamics (SGLD) and the stochastic gradient Hamiltonian Monte Carlo (SGHMC), can scale to large datasets for practical applications.
SGLD. The objective of MCMC is to obtain random samples of the posterior distribution of model parameters (with the unnormalized form) $p(\mathbf{w}|\mathcal{D})\propto\exp(-U(\mathbf{w}))$, where $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})$ is the potential energy function. The idea of dynamics-based sampling for $p(\mathbf{w}|\mathcal{D})$ is to simulate a continuous-time stochastic process whose stationary distribution is the target posterior. Combining Langevin dynamics (Neal et al., 2011) with stochastic Robbins-Monro optimization algorithm (Robbins & Monro, 1951), Welling & Teh (2011) proposed SGLD for scalable MCMC algorithms. The approximated discretization transition rule for posterior samples is
@@
where $\epsilon_{t}$ is the step-size, $\eta_{t}$ is a standard Gaussian, and $\nabla\tilde{U}(\mathbf{w})$ uses mini-batch stochastic gradient for the likelihood term as $\nabla\tilde{U}(\mathbf{w})=-\frac{N}{n}\sum_{i=1}^{n}\nabla\log p(y_{i}|x_{i},\mathbf{w})-\nabla\log p(\mathbf{w})$ and $n\ll N$. With the decreasing step-size $\epsilon_{t}$, the stochastic gradient noise tends to zero faster than the injected Gaussian noise $\eta_{t}$, and the discretization error is eliminated, then the MH step can be ignored.
SGHMC. To enable a more efficient exploration in the state space, Hamiltonian dynamics (Duane et al., 1987; Neal, 2012) introduces a momentum variable $\mathbf{z}\in\mathbb{R}^{d}$, and the new joint distribution is $p(\mathbf{w}|\mathcal{D})p(\mathbf{z})\propto\exp(-H(\mathbf{w},\mathbf{z}))$ with corresponding energy function $H(\mathbf{w},\mathbf{z})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})-\log p(\mathbf{z})$, where $p(\mathbf{z})$ is commonly assumed as a zero-mean Gaussian $\mathcal{N}(\mathbf{z}|0,M)$. The sample rule in the full-batch Hamiltonian Monte Carlo is defined as following with an MH step (Neal, 2012; Cobb & Jalaian, 2021):
@@
Like SGLD, SGHMC uses stochastic gradient $\nabla\tilde{U}(\mathbf{w})$ for the sampling of momentum $\mathbf{z}$, also with an additional friction term $\epsilon_{t}M^{-1}\mathbf{z}_{t}$ and a Gaussian injected noise to guarantee the (marginal) stationary distribution of the Hamiltonian dynamics still be the target posterior distribution (Chen et al., 2014). The specific update rule for samples of $\mathbf{z}$ is as: $\mathbf{z}_{t+1}=\mathbf{z}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})-\epsilon_{t}CM^{-1}\mathbf{z}_{t}+\sqrt{2C\epsilon_{t}}\eta_{t}$, where $C>0$ is a scalar coefficient.
3 Functional Stochastic Gradient MCMC
@@
Generally, dynamics-based MCMC approaches are derived from the Itô diffusion (Øksendal, 2003), which is similar to the Brownian motion of the particle subjected to a potential in fluid physics. In this section, to solve the parameter-space prior issues in existing SGMCMC methods, we propose a new functional SGMCMC scheme via a newly designed diffusion dynamics, which is capable of incorporating more informative functional priors. In Section 3.1, we first review the Itô diffusion for dynamics-based MCMC. Then we design a new functional Langevin dynamics directly for the function mapping defined by the BNNs. We prove that the stationary distribution of such functional Langevin dynamics is the target posterior distribution over functions. Furthermore, in Section 3.2, we design a functional Hamiltonian dynamics and the corresponding stochastic gradient sampling scheme for BNNs. We also prove it can guarantee the desired invariant distribution.
3.1 Functional SGLD for BNNs
An Itô diffusion in n-dimensional Euclidean space is defined as a stochastic differential equation (SDE) to describe the evolution of a continuous-time stochastic process $X:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}$ on a probability space $(\Omega,\Sigma,P)$. The general form of an Itô diffusion driven by the standard Wiener process (Brownian motion) is as follows:
$\mathrm{d}X_{t}=\mu(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},$ (3)
where $X_{t}$ represents the state of the stochastic process at time $t$, $B\in\mathbb{R}^{n}$ is an n-dimensional Brownian motion, $\mu(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a vector field denoting the deterministic drift term and matrix field $\sigma(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ is known as the diffusion coefficient for $X_{t}$. Both $\mu(\cdot)$ and $\sigma(\cdot)$ are assumed to satisfy the usual Lipschitz continuity condition (Ghosh, 2010). And $\mathrm{d}B_{t}$ is the increment of a Wiener process distributed as $\mathrm{d}B_{t}\sim\mathcal{N}(0,\mathrm{d}t\cdot\mathbf{I}_{n})$. Let $p(x,t)$ be the probability density of random variable $X_{t}$, and the corresponding evolution of this probability density of the diffusion process can be described by the Fokker–Planck equation (Risken, 1996), which is the key to derive the dynamics-based MCMC (see Appendix B).
@@ 3.1 Functional SGLD for BNNs
Naturally, the idea of existing dynamics-based SGMCMC methods for BNNs is to simulate a continuous diffusion of network weight whose stationary distribution is the target posterior $p(\mathbf{w}|\mathcal{D})$ with appropriately designed drift term and diffusion coefficient. To incorporate informative functional prior into the posterior inference process, we now directly consider the dynamics of function mapping $f(\cdot;\mathbf{w})$ defined by BNNs. According to Itô Lemma (Itô, 1951), $f(\cdot;\mathbf{w})$ is itself an Itô diffusion (Brzeźniak et al., 2008) (see Appendix A). That is, for exiting SGMCMC methods, the stochastic process $X$ simulated in Itô diffusion is usually for network weight or its augmentation, which we now lift to the stochastic function mapping of BNNs directly as
$\mathrm{d}f_{t}=\mu(f_{t})\mathrm{d}t+\sigma(f_{t})\mathrm{d}B_{t}.$ (4)
In a probability space $(\Omega,\mathcal{F},P)$ (separable and complete Polish space), suppose $p(f)$ is the prior distribution over functions, such as the classical Gaussian process (GP) prior that can easily encode prior knowledge about function properties. $p(\mathbf{Y_{\mathcal{D}}}|f(\mathbf{X_{\mathcal{D}}}))$ is the likelihood function with observation $\mathcal{D}=\left\{\mathbf{X_{\mathcal{D}}},\mathbf{Y_{\mathcal{D}}}\right\}$ and $p(f|\mathcal{D})$ is the posterior over functions. For brevity, we abbreviate them as $p(f)$, $p(\mathcal{D}|f)$ and $p(f|\mathcal{D})$, respectively. Let $U(f)=-\log p(\mathcal{D}|f)-\log\frac{\mathrm{d}p}{\mathrm{d}r_{\mathcal{H,F}}}(f))$ be the potential energy functional, where $r_{\mathcal{H,F}}$ is the corresponding Hausdorff / Riemannian measure, and $\mu(f)=-\nabla_{f}U(f),\sigma(f)=\sqrt{2}\cdot\mathbf{I}_{\infty}$ for the function mapping $f$ of BNNs, we have the following functional Langevin dynamics:
$\mathrm{d}f_{t}=-\nabla_{f}U(f_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$ (5)
and it is easy to verify that the stationary distribution for $p(f,t)$ is exactly the target posterior distribution over functions as $\pi(f)=\exp(-U(f))=p(f|\mathcal{D})$.
However, even for the finite input case, it is intractable to discretize and sample the above diffusion of function mapping in practice directly. Considering that a function mapping $f(\cdot;\mathbf{w})$ of BNNs is completely determined by the network weights $\mathbf{w}$ (assuming the network architecture is fixed), we solve this issue by transforming the functional Langevin dynamics to a weight-space dynamics of which stationary distribution is still $p(f|\mathcal{D})$. We firstly design the following specific Riemannian functional Langevin dynamics for $f(\cdot;\mathbf{w})$ as follows:
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=$ $\mu(f_{t}(\cdot;\mathbf{w}))\mathrm{d}t+\sigma(f_{t}(\cdot;\mathbf{w}))\mathrm{d}B_{t}$ (6)
$=$ $\left[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f\right]\mathrm{d}t$
$+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t},$
@@
where $\nabla_{\mathbf{w}}f$ is the gradient of $f(\cdot;\mathbf{w})$ w.r.t. weight $\mathbf{w}$, $H_{\mathbf{w}}f$ is the Hessian matrix of $f(\cdot;\mathbf{w})$ w.r.t. $\mathbf{w}$. The stationary distribution of this dynamics is guaranteed to be $p(f|\mathcal{D})$ by the following Proposition.
@@
The stationary distribution of the functional Langevin dynamics defined in Equation 6 is the target posterior distribution over functions $p(f|\mathcal{D})$.
@@
The proof is given in Appendix B. ∎
Secondly, we achieve the function-space to weight-space transformation based on the relationship between the dynamics of a random variable with the dynamics of its function. According to the Itô Lemma (Itô, 1951), given the general diffusion framework for weights as $\mathrm{d}\mathbf{w}_{t}=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$, the differential of the function mapping $f(\cdot;\mathbf{w})$ can be denoted as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ (7)
$=(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}\mathbf{w}_{t}+\frac{1}{2}(\mathrm{d}\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\mathrm{d}\mathbf{w}_{t}$

### v2

@@ Abstract
Classical parameter-space Bayesian inference for Bayesian neural networks (BNNs) suffers from several unresolved prior issues, such as knowledge encoding intractability and pathological behaviours in deep networks, which can lead to improper posterior inference. To address these issues, functional Bayesian inference has recently been proposed leveraging functional priors, such as the emerging functional variational inference. In addition to variational methods, stochastic gradient Markov Chain Monte Carlo (MCMC) is another scalable and effective inference method for BNNs to asymptotically generate samples from the true posterior by simulating continuous dynamics. However, existing MCMC methods perform solely in parameter space and inherit the unresolved prior issues, while extending these dynamics to function space is a non-trivial undertaking. In this paper, we introduce novel functional MCMC schemes, including stochastic gradient versions, based on newly designed diffusion dynamics that can incorporate more informative functional priors. Moreover, we prove that the stationary measure of these functional dynamics is the target posterior over functions. Our functional MCMC schemes demonstrate improved performance in both predictive accuracy and uncertainty quantification on several tasks compared to naive parameter-space MCMC and functional variational inference.
@@ 1 Introduction
Compared to traditional deep neural networks that rely on a single pattern of model parameters only, Bayesian neural networks (BNNs) demonstrate principled predictive uncertainty estimation and improved generalization by integrating models under the posterior distribution over network parameters (Neal,, 1995; Gal et al.,, 2016; Wilson and Izmailov,, 2020). As Bayesian inference methods continue to advance and gain success within the deep learning community, BNNs have been applied in a wide variety of domains and tasks, including dissolution prediction of planetary systems (Cranmer et al.,, 2021), medical diagnosis as diabetic detection (Filos et al.,, 2019; Band et al.,, 2022), classification of radio galaxies (Mohan and Scaife,, 2024), and other safety-critical environments (Rudner et al.,, 2022).
@@
Choosing an appropriate prior is critical for BNNs. The common practice is to use isotropy Gaussian priors for all network parameters, often referred to as parameter-space priors. However, these parameter-space priors have shown several problematic issues. For instance, as the depth of the network increases, the function samples of Gaussian prior over parameters tend to be horizontal (Duvenaud et al.,, 2014; Matthews et al.,, 2018). Additionally, with ReLU activations, the prior distribution of unit outputs becomes more heavy-tailed for deeper architectures.(Vladimirova et al.,, 2019; Tran et al.,, 2022). Furthermore, the effects of this parameter-space prior to posterior inference remain an open question due to the non-interpretability of network parameters and the complexity of network architectures Fortuin et al., (2021); Wild et al., (2022), which make it challenging to translate valid prior knowledge into a corresponding prior distribution over network parameters. Due to these unresolved issues with parameter-space priors, there has been growing interest in performing Bayesian inference directly in function space, where more informative stochastic processes can serve as priors. For example, Gaussian processes (GPs), which can easily encode prior knowledge about the potential functions via kernel functions, have been widely used as priors in functional variational inference for BNNs (Sun et al.,, 2019; Ma and Hernández-Lobato,, 2021; Rudner et al.,, 2022; Pielok et al.,, 2022). These functional variational inference approaches have shown improved performance compared to parameter-space variational inference.
@@
For posterior inference given a prior, Markov Chain Monte Carlo (MCMC) is another valid method for BNNs apart from variational inference. Unlike variational inference, which makes strong assumptions about the posterior distribution, non-parametric MCMC methods are asymptotically exact, making them the gold standard for Bayesian inference (Alexos et al.,, 2022). Even better, the computationally expensive MCMC methods became feasible for modern BNNs with large data following the introduction of their scalable stochastic gradient variants (SGMCMC) (Welling and Teh,, 2011; Patterson and Teh,, 2013; Chen et al.,, 2014; Zhang et al.,, 2020; Nemeth and Fearnhead,, 2021). However, existing MCMC methods for BNNs typically perform in parameter space and are based on the diffusion dynamics designed in terms of the posterior over parameters, inheriting similar prior issues as parameter-space variational inference. Extending these methods to function space, however, is a non-trivial task. It requires the design of novel dynamics that not only sample effectively from the function space but also guarantee that the resulting stationary distribution corresponds to the target posterior over functions. This involves addressing the inherent complexities of functional spaces while ensuring the correctness and feasibility of the sampling process.
@@
In this work, we propose novel functional MCMC schemes grounded in the fundamental Itô lemma for stochastic calculus (Itô,, 1951; Brzeźniak et al.,, 2008; Øksendal,, 2013). Specifically, the dynamics of BNNs in function space are designed by introducing new potential energy functionals that can incorporate functional priors, such as GPs, into the diffusion processes and simultaneously guarantee the target posterior over functions as the stationary measure. Our main contributions are as follows:
  • We propose novel functional diffusion dynamics, along with their stochastic gradient versions for BNNs that effectively incorporate more informative functional priors into posterior inference. Additionally, we derive corresponding tractable MCMC samplers in parameter space, making them computationally feasible while maintaining the expressiveness of functional priors.
@@
  • We prove that the stationary measure would be the target posterior over functions under our designed functional Langevin dynamics and functional Hamiltonian dynamics.
@@
  • Our methods demonstrate improved predictive performance and uncertainty estimation in several benchmark tasks compared to parameter-space SGMCMC and parameter-/function-space variational inference methods.
@@ 2.1 Bayesian Neural Networks
Bayesian neural networks (BNNs) place probability distribution over network parameters rather than single fixed values used in standard neural networks. Given a dataset $\mathcal{D}=\left\{x_{i},y_{i}\right\}_{i=1}^{N}=\left\{\mathbf{X_{\mathcal{D}}},\mathbf{Y_{\mathcal{D}}}\right\}$ , where $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ are the training inputs and $y_{i}\in\mathcal{Y}\subseteq\mathbb{R}^{c}$ are the corresponding targets, BNNs are stochastic neural networks characterized by random network parameters $\mathbf{w}\in\mathbb{R}^{k}$, treated as a k-dimensional multivariate random variable defined on a probability space $(\Omega,\mathcal{A},P)$. The prior distribution over parameters is denoted by $p_{0}(\mathbf{w})$, and the likelihood evaluated on the training data is represented by the conditional distribution $p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})$. The posterior over parameters can then be inferred using Bayes’ theorem, yielding the canonical form $p(\mathbf{w}|\mathcal{D})\propto p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})p_{0}(\mathbf{w})$. Furthermore, the predictive distribution for test data $\left\{x^{*},y^{*}\right\}$ is given by the expectation under the posterior as $p(y^{*}|x^{*},\mathcal{D})=\int p(y^{*}|x^{*},\mathbf{w})p(\mathbf{w}|\mathcal{D})\mathrm{d}\mathbf{w}$. All notations used in this paper are listed in Table 5 in Appendix A.
@@ 2.2 Stochastic Gradient MCMC for BNNs
Dynamics-based sampling methods that leverage the gradients of the target log posterior can offer an efficient exploration of the parameter space (Ma et al.,, 2015; Izmailov et al.,, 2021; Li and Zhang,, 2024). Furthermore, their stochastic gradient variants, such as the stochastic gradient Langevin dynamics (SGLD) (Welling and Teh,, 2011) and the stochastic gradient Hamiltonian Monte Carlo (SGHMC) (Chen et al.,, 2014; Baker et al.,, 2019), are capable of scaling to large datasets for practical applications.
SGLD.
The objective of MCMC for BNNs is to generate random samples from the posterior distribution over parameters (with the unnormalized form) as $p(\mathbf{w}|\mathcal{D})\propto\exp(-U(\mathbf{w}))$, where $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})-\log p_{0}(\mathbf{w})$ is the potential energy function. By combining Langevin dynamics (Neal et al.,, 2011) with the stochastic Robbins-Monro optimization algorithm (Robbins and Monro,, 1951), Welling and Teh, (2011) proposed SGLD as a scalable MCMC algorithm. The approximated discretization transition rule for posterior samples is
@@
where $\epsilon_{t}$ represents the step-size, and $\eta_{t}$ is a standard Gaussian noise term. The gradient $\nabla\tilde{U}(\mathbf{w})$ is computed using a mini-batch stochastic gradient for the likelihood term given by $\nabla\tilde{U}(\mathbf{w})=-\frac{N}{n}\sum_{i=1}^{n}\nabla\log p(y_{i}|x_{i};\mathbf{w})-\nabla\log p(\mathbf{w})$ with $n\ll N$. As the step-size $\epsilon_{t}\rightarrow 0$, the stochastic gradient noise vanishes more quickly than the injected Gaussian noise $\eta_{t}$, which ensures the discretization error is eliminated. Hence, the Metropolis-Hastings (MH) correction step can be ignored.
SGHMC.
To enable more efficient exploration of the state space, Hamiltonian dynamics (Duane et al.,, 1987; Neal,, 2012) introduces a momentum variable $\mathbf{z}\in\mathbb{R}^{k}$. The new target joint distribution is $p(\mathbf{w}|\mathcal{D})p(\mathbf{z})\propto\exp(-H(\mathbf{w},\mathbf{z}))$ with corresponding energy function $H(\mathbf{w},\mathbf{z})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})-\log p(\mathbf{z})$, where $p(\mathbf{z})$ is commonly assumed to be a zero-mean Gaussian $\mathcal{N}(\mathbf{z}|0,M)$. The sampling rule in the full-batch Hamiltonian Monte Carlo is defined as following with an MH step (Neal,, 2012; Cobb and Jalaian,, 2021):
@@
Like SGLD, SGHMC uses stochastic gradient $\nabla\tilde{U}(\mathbf{w})$ for the sampling of momentum $\mathbf{z}$, but with an additional friction term $\epsilon_{t}M^{-1}\mathbf{z}_{t}$ and a Gaussian injected noise to ensure that the (marginal) stationary distribution of the Hamiltonian dynamics remains the target posterior distribution (Chen et al.,, 2014). The specific update rule for the momentum $\mathbf{z}$ is given by $\mathbf{z}_{t+1}=\mathbf{z}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})-\epsilon_{t}CM^{-1}\mathbf{z}_{t}+\sqrt{2C\epsilon_{t}}\eta_{t}$, where $C>0$ is a scalar coefficient.
3 Functional Dynamics-based MCMC
@@
In this section, we address the prior issues present in parameter-space MCMC by proposing two novel functional MCMC schemes based on newly designed diffusion dynamics, which allow for the incorporation of more informative functional priors. In Section 3.1, we introduce novel functional Langevin dynamics designed directly for the random function mapping defined by BNNs, along with a corresponding tractable stochastic gradient MCMC sampler for network parameters. We demonstrate that the stationary probability measure of this functional Langevin dynamics corresponds to the target posterior over functions. Additionally, in Section 3.2, we develop functional Hamiltonian dynamics also with a stochastic gradient version for BNNs. We prove that this scheme also guarantees the desired functional posterior as the stationary measure.
3.1 Functional Langevin Dynamics for BNNs
@@
Dynamics-based MCMC techniques are grounded in the general framework of Itô diffusion (Øksendal,, 2003). The main idea behind parameter-space MCMC methods for BNNs is to simulate a continuous diffusion process of network parameters $\mathbf{w}$ such that its stationary distribution is the target posterior $p(\mathbf{w}|\mathcal{D})$, using appropriately designed drift term and diffusion coefficient (see Appendix B). Naturally, to incorporate more informative functional prior into the posterior inference process, we now consider lifting the parameter-space Langevin dynamics onto function space.
Suppose $f(\cdot;\mathbf{w}):\mathcal{X}\times\mathbb{R}^{k}\rightarrow\mathcal{Y}$ is the product-measurable random function mapping defined by a BNN. Assums $f(\cdot;\mathbf{w})\in\mathbb{H}$, where $\mathbb{H}$ is an infinite-dimensional function (Polish) space with Borel $\sigma$-algebra $\mathcal{B}(\mathbb{H})$. The likelihood is defined as $p$: $\mathcal{Y}\times\mathbb{H}\rightarrow[0,\infty)$, mapping $(\mathcal{Y},f(\mathcal{X};\mathbf{w}))\mapsto p(\mathcal{Y}|f(\mathcal{X};\mathbf{w}))$, where $\mathcal{Y}\subseteq\mathbb{R}^{c}$ is Borel measurable and $p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$ denotes the likelihood evaluated on the training data. Let $\mathcal{P}(\mathbb{H})$ represent the space of Borel probability measures on $\mathcal{B}(\mathbb{H})$, and assume the functional prior measure $P_{0}\in\mathcal{P}(\mathbb{H})$ with topological support $\operatorname{supp}(P_{0})$, such as the classical GP prior denoted as $P_{0}\sim\mathcal{GP}(\mathbf{m},\mathbf{K})$. The posterior measure $P_{f|\mathcal{D}}\in\mathcal{P}(\mathbb{H})$, induced by the functional prior and likelihood, is defined by the Radon–Nikodym derivative as $P_{f|\mathcal{D}}(\mathrm{d}f)\propto\exp(-\Phi(f))P_{0}(\mathrm{d}f)$, where $\Phi(f)$ is the negative log-likelihood as $-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$ (Matthews et al.,, 2016; Lambley,, 2023).
According to Itô Lemma (Itô,, 1951), $f(\cdot;\mathbf{w})$ itself follows an Itô diffusion (Brzeźniak et al.,, 2008) (see Appendix B). Thus, we propose to directly simulate the diffusion process of the random function mapping $f(\cdot;\mathbf{w})$ of BNNs, where one can incorporate informative functional prior, $P_{0}$. Suppose $\varnothing\neq E\subseteq\operatorname{supp}(P_{0})$, and let $I_{0}:E\rightarrow\mathbb{R}$ be an Onsager–Machlup (OM) functional for $P_{0}$, which can be heuristically interpreted as the negative logarithm of $P_{0}$ (Ayanbayev et al.,, 2021) (see Appendix B). Similar to parameter-space Langevin dynamics, the Langevin dynamics in function space can be generally represented by the stochastic differential equation (SDE): $\mathrm{d}f_{t}=-\nabla U(f_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}$, where $B$ is a standard Wiener process ((Brownian motion), and potential energy functional $U:E\rightarrow\mathbb{R}$ is given by $U(f)=\Phi(f)+I_{0}(f)$. This functional $U(f)$ is proven to be an OM functional for $P_{f|\mathcal{D}}$ (Lambley,, 2023), ensuring that the target posterior measure $P_{f|\mathcal{D}}$ as the stationary measure $\pi(f)$. However, discretizing and sampling from this diffusion process in function space is intractable due to the infinite-dimensional nature of the problem.
Since $f(\cdot;\mathbf{w})$ is completely determined by the network parameters $\mathbf{w}$ (given a fixed network architecture), we address this challenge by designing a specific functional Langevin dynamics for $f(\cdot;\mathbf{w})$, which maintains the stationary measure $P_{f|\mathcal{D}}$. We then transform this into the corresponding parameter-space dynamics for network parameters. Suppose that function space $\mathbb{H}$ is equipped with a Riemannian manifold structure (Wang et al.,, 2022), we first design the following functional Riemannian Langevin dynamics for $f(\cdot;\mathbf{w})$ on $\mathbb{H}$:
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=\mu(f_{t}(\cdot;\mathbf{w}))\mathrm{d}t+\sqrt{2G^{-1}(f_{t}(\cdot;\mathbf{w}))}\mathrm{d}B_{t}$ (3)
$=\Big{[}-(\nabla_{\mathbf{w}}f_{t})^{T}(\nabla_{\mathbf{w}}f_{t})\nabla_{f}(-\log p(\mathbf{Y}_{\mathcal{D}}|f_{t}(\mathbf{X}_{\mathcal{D}};\mathbf{w}))+$
$\quad I_{0}(f_{t}))+H_{\mathbf{w}}f_{t}\Big{]}\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f_{t})^{T}\mathrm{d}B_{t},$
@@
where $\nabla_{\mathbf{w}}f$ denotes the gradient of $f(\cdot;\mathbf{w})$ w.r.t. parameters $\mathbf{w}$, $H_{\mathbf{w}}f$ is the Hessian matrix, and $G^{-1}(f(\cdot;\mathbf{w}))=(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)$. The stationary probability measure of this dynamics is guaranteed to be $P_{f|\mathcal{D}}$ as stated in the following proposition.
@@
The stationary probability measure of the functional Langevin dynamics defined in Equation 3 is the target posterior over functions $P_{f|\mathcal{D}}$.
@@
The proof is given in Appendix C. ∎
To establish a tractable functional MCMC scheme, we then transform the proposed functional Riemannian Langevin dynamics for $f(\cdot;\mathbf{w})$ into the corresponding Langevin dynamics for the network parameters $\mathbf{w}$. Note that the differential of the function mapping $f(\cdot;\mathbf{w})$ is given by the Itô Lemma (Itô,, 1951) as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}\mathbf{w}_{t}+\frac{1}{2}(\mathrm{d}\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\mathrm{d}\mathbf{w}_{t}$ (4)